OpenStudy (anonymous):

Can you guys give me a hint on how to answer this? "A ball, rolling at 3m/s, rolls of a dock and hits the water 0.3m from the dock. How high is the dock above the water?"

OpenStudy (anonymous):

I would use the equation 1/2(mass)(velocity)^2=(mass)(gravity)(height) aka Kinetic Energy=Gravitational Potential Energy Cancel out the masses on either side of the equation so 1/2v^2=gh and plug in the values for velocity and gravity. Then solve for h. Hope this helps

OpenStudy (anonymous):

The above approach ^ will not work, as the velocity given is in a horizontal direction. To then relate it to gravity would require it to be in the vertical direction. This problem can be solved by simply considering how long the ball would need to travel if going 3 m/s to have a displacement of 0.3 m. Then you can find out how high the dock needs to be in order for the ball to fall to the water surface in the time you just found. In summary: (1) Solve for time taken for v = 3 m/s to displace x = 0.3 m (2) Use time 't' from (1) to find height needed for free-falling object using: h = 0.5*g*t^2

OpenStudy (anonymous):

I see what you are saying. I don't know what I was thinking with my previous answer! Anyways, I believe the approach you are using would work. For t from (1), I think the equation distance=velocity*time would work, but I'm not quite sure since I can't remember if its okay to use d=vt when velocity is constant.

OpenStudy (anonymous):

Thank you!